# Lifshitz Quantum Mechanics and Anisotropic Josephson Junction

### v1

@@ Abstract
We consider quantum mechanics in spacetime with time anisotropy. Such Lifshitz quantum mechanics is characterized by a kinetic term with fractional derivatives. We show that, contrary to popular belief, local conservation of probability is respected when the probability current is properly identified. As an application we consider a Josephson Junction with an anisotropic insulating layer. We show that anisotropy affects the tunneling rate and can greatly enhance the working of Josephson Junction.
@@ I Introduction
Many interesting systems in physics are anisotropic. For example, liquid crystal, magnetic materials and critical systems in condensed matter physics are ubiquitously anisotropic. The cosmic microwave background also display traces of anisotropy that are thought to have originated from quantum fluctuations occurred during inflation. Often the fundamental physical law is isotropic but the presence of a background leads to anisotropy. An interesting instant of anisotropy is when physics is invariant under the Lifshitz scaling transformation
@@
Here spatial isotropy is maintained while time is anisotropic for an anisotropy exponent $z\neq 1$. Such Lifshitz scaling symmetry occurs naturally as symmetry of the critical point of renormalization group (RG) flow for anisotropic systems. Lifshitz scale invariant system can be recognized easily by the dispersion relation of its massless modes:
@@
where one can extract not only the anisotropy exponent $z$ but also some mass scale that is associated with the anisotropy. As the only possible form that is compatible with the scaling symmetry (I.1), (I.2) is independent of the underlying origin of the anisotropy. The symmetry for $z=2$ was first proposed by Lifshitz [1] to describe the tricritical point of liquid helium [2]. The dispersion relation (I.2) is also featured in modified gravity scenario with a ghost condensate [3]. In condensed matter physics, a well known example is the Rokhsar-Kivelson (RK) Quantum dimer model [4, 5] where a continuum $z=2$ Lifshitz scalar field theory emerges at the critical point.
The $z=2$ anisotropic scaling symmetry (I.1) has been intensively studied from the field theory point of view [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. For example, it is known that Lifshitz scalar field theory admits ground state called the RK vacuum, with the properties that it is local and is given by a superposition of quantum states determined by some classical probability distribution [4]. The point $z=2$ is somewhat special since the Lifshitz scaling transformation (I.1) combines with the Galilean boost to form the Schrodinger group only for this value of $z$. Study of field theory with general $z$ was initiated in [20] where with the help of fractional calculus. It was shown that, independent of whether there is an enlarged Schrodinger symmetry or not, the ground state of the Lifshitz scalar field always takes the form of RK. The entropic properties of the RK vacuum were also studied, both in field theory and by holography, in [20].
Clearly, the physics of anisotropy is a subject of its own interest that goes beyond quantum field theory. With general applicability in mind, we explore in this paper the physical effect of anisotropy in quantum mechanics. In particular we consider a kind of time anisotropy of the background space time specified by (I.1). With the dispersion relation fixed to be of the form (I.1), quantum mechanics of free particle in the anisotropic spacetime reads
@@
where $H_{0}$ is the free Hamiltonian, $m$ is the mass of the particle and $M$ is a mass scale associated with the anisotropy. The fractional Laplacian $\Delta_{z}$ of order $z$ is defined by its action on the basis of exponential functions by
$\triangle_{z}\,e^{{\bf k}\cdot{\bf x}}=({\bf k}^{2})^{z/2}e^{{\bf k}\cdot{\bf x}}.$ (I.4)
The free (I.3) is Lifshitz scale invariant. One may also add a potential and consider the Lifshitz quantum mechanics defined by
@@
Note that in general the scaling symmetry is broken except for some special form of the potential, e.g. $V\sim 1/|x|^{z}$, in which case we obtain the Lifshitz invariant quantum mechanics. In analogous to conformal quantum mechanics [21], it may be interesting to study further this kind of Lifshitz invariant quantum mechanics.
Here is a summary of the paper. In section 2, we consider the Lifshitz quantum mechanics in general. There is a belief in the literature that local conservation of probability is violated by a source term in Lifshitz quantum mechanics. It turns out it is in fact the probability current that is modified by anisotropy. And once the correct probability current is employed, the local conservation of probability is respected. In section 3, we consider an anisotropic Josephson Junction that is made up of an anisotropic insulating layer sandwiched by two normal superconducting materials. We show that the critical current can be greatly enhanced by anisotropy. We argue that anisotropic Josephson Junction has great potential in quantum computing.
@@ II Lifshitz Quantum Mechanics
Consider Lifshitz quantum mechanics (I.5). This kind of quantum mechanics is also refereed to as fractional quantum mechanics, see for example [22] for some discussion. It is easy to show that the Lifshitz quantum mechanics (I.5) is unitary and admits the local conservation law
@@
where the probability density takes the usual form $\rho=\psi^{*}\psi$, while the probability current density is modified by anisotropy. As the fractional derivative is defined using the basis of exponential functions, $\pmb{J}$ can be most easily obtained in the momentum space and we have
@@
This result can be obtained from the Noether theorem associated with the phase symmetry $\psi\to\psi e^{i\theta}$ of quantum mechanics or by the direct manipulation of the wave equation (I.5). Equivalently in coordinate space, it is
@@
$\mathbf{J}=\lambda_{M}^{z-2}\frac{\hbar}{2mi}\frac{\nabla_{x}-\nabla_{x^{\prime}}}{\nabla_{x}^{2}-\nabla_{x^{\prime}}^{2}}\left[(\nabla_{x}^{2})^{z/2}-(\nabla_{x^{\prime}}^{2})^{z/2}\right]\psi^{*}(\mathbf{x}^{\prime},t)\psi(\mathbf{x},t)\Big{|}_{\mathbf{x}^{\prime}=\mathbf{x}},$ (II.4)
where $\lambda_{M}:=\hbar/M$ is a length scale associated with the anisotropy.
We remark that in the form (II.2), it is easy to see that instead of using the kernel $\frac{{\bf q}+{\bf p}}{{\bf q}^{2}-{\bf p}^{2}}$ in (II.2), one may also use any vector $\bm{\beta}$ that satisfies
$\bm{\beta}\cdot({\bf q}-{\bf p})=1$ (II.5)
@@
to produce a $\pmb{J}$ that observes the conservation law. For example, $\bm{\beta}=\frac{{\bf q}-{\bf p}}{|{\bf q}-{\bf p}|^{2}}$ is good too. To see this, suppose $\bm{\beta}_{2},\bm{\beta}_{1}$ both satisfy (II.5), then $(\bm{\beta}_{2}-\bm{\beta}_{1})$ is perpendicular to $({\bf q}-{\bf p})$, i.e. $\bm{\beta}_{2}-\bm{\beta}_{1}=({\bf q}-{\bf p})\times{\bf s}$ for some ${\bf s}$. This implies that the corresponding $\pmb{J}$’s differ from each other by a curl term, and hence both satisfy the same conservation law. This kind of ambiguity is always there even for the standard quantum mechanics although it is not usually mentioned. Like the Poynting vector, additional input from the physical setting such as symmetry, boundary condition or a good understanding of the underlying dynamics, is usually needed to fix the choice of $\pmb{J}$ properly. This is the case for the example that we will consider below. We remark also that the local conservation of probability in fractional quantum mechanics has been discussed in [23, 24, 25] where a source term has been found, and which has led to the confusion of non-conservation of local probability. As we showed above, the probability current is indeed modified by anisotropy and local conservation of probability is observed in fractional quantum mechanics. This conclusion is important as we will use below the local conservation of probability to derive a statement about the macroscopic electric current in the Josephson Junction system.
@@ III Anisotropic Josephson Junction
Suppose the Cooper pairs have typical energy $E$ in the regions 1, 2 and the insulator is represented as a potential barrier $V_{0}>E$ (see figure 2). Then the wave equation inside the insulator,
$\lambda_{M}^{z-2}\frac{\hbar^{2}\Delta_{z}\psi}{2m}=(V_{0}-E)\psi$ (III.2)
has solution
@@
Here $\zeta$ is the tunneling penetration length and $\zeta_{0}$ is the tunneling penetration length in the undeformed $z=2$ (isotropic) case. Note that the penetration length is modified by anisotropy through the presence of the anisotropy length scale $\lambda_{M}$ and anisotropy index $z$. Next let us determine the tunneling current.
@@
Note that here we have assumed the Junction to be uniform in the transverse directions and so we have effectively a one spatial dimension system. In this case there is no room for any curl term and the current is uniquely defined by (II.4). We obtain
@@
The coefficients $C_{1},C_{2}$ can be determined by imposing the boundary conditions at $x=\pm a$:
@@
Taking into account the Cooper pair has charge $q=2e$, we obtain the first Josephson equation for the tunneling current
$J=J_{c}\sin\varphi,\quad\varphi=\theta_{2}-\theta_{1}.$ (III.7)
Here the critical current $J_{c}$ is given by
$J_{c}=\gamma J_{c0}$ (III.8)
@@
$J_{c0}=\frac{q\hbar\sqrt{n_{1}n_{2}}}{m\zeta_{0}}\frac{1}{\sinh 2a/\zeta_{0}},\quad\gamma=\frac{z\zeta}{2\zeta_{0}}\frac{\sinh 2a/\zeta_{0}}{\sinh 2a/\zeta},$ (III.9)
where $J_{c0}$ is the critical current in the isotropic case and $\gamma$ is an modification factor due to anisotropy. If we consider the application of a voltage $V$ across the Josephson Junction, then the phase becomes time dependent and we have the second Josephson equation,
@@
[…]
@@
Note that in the usual isotropic case, the critical current $J_{c0}$ is exponentially suppressed by a factor of $e^{-2a/\zeta}$ since the Cooper pairs cannot tunnel deep enough into the insulating region. However this may change in the anisotropic situation. In fact, the Josephson current can be enhanced ($\gamma>1$) if the penetration length is made bigger ($\zeta>\zeta_{0}$). This happens if $z<2$ and $\lambda_{M}<\zeta_{0}$ or $z>2$ and $\lambda_{M}>\zeta_{0}$. Thus an anisotropic material has the potential of enhancing the working of Josephson Junction. In particular, if the anisotropic insulating material can be engineered such that the penetration length $\zeta$ get to become of the same order as $a$, then tunneling can become much more effective and $J_{c}$ may be greatly enhanced. As illustration, let us consider an isotropic sample with $\zeta_{0}/2a=1/10$, in this case the critical current is suppressed with a factor of $\sim e^{-10}\sim 10^{-5}$. Introducing anisotropy, for example with $z=0.8$ and $\zeta_{0}/\lambda_{M}=10$, we can have a huge amplification factor $\gamma\sim 10^{5}$; or with $z=5$ and $\zeta_{0}/\lambda_{M}=1/10$, we have $\gamma\sim 10^{4}$. We remark that the dispersion relation (I.2) is acausal for $z<1$ due to the existence of superluminal modes [26] and the violation of the null energy conditions of the holographic dual [27]. Therefore a junction with $z\geq 1$ may be more physical. It is interesting to verify this prediction of enhanced tunneling experimentally.

### v2

@@ Abstract
We consider quantum mechanics in spacetime with anisotropy in time. Such Lifshitz quantum mechanics is characterized by a kinetic term with fractional derivatives. We show that, contrary to a common claim in the literature, local conservation of probability is respected when the probability current is properly identified. As an application we consider a Josephson Junction with an insulating layer exhibiting Lifshitz anisotropy. We show that anisotropy modifies the tunneling rate and can significantly enhance the performance of the Josephson Junction.
@@ I Introduction
Many interesting systems in physics are anisotropic. For example, liquid crystals, magnetic materials and critical systems in condensed matter physics are ubiquitously anisotropic. The cosmic microwave background also display traces of anisotropy that are thought to have originated from quantum fluctuations occurred during inflation. Often the fundamental physical laws are isotropic but the background structure induces anisotropy. An interesting instance of anisotropy is when physics is invariant under the Lifshitz scaling transformation
@@
Here spatial isotropy is maintained while time is anisotropic for a dynamical exponent $z\neq 1$. Such Lifshitz scaling symmetry occurs naturally as the symmetry at the critical point of renormalization group (RG) flow for anisotropic systems. Lifshitz scale invariant system can be recognized easily by the dispersion relation of their massless modes:
@@
from which one can extract not only the dynamical exponent $z$ but also a mass scale associated with the anisotropy. As the only possible form that is compatible with the scaling symmetry (I.1), (I.2) is independent of the underlying origin of the anisotropy. The symmetry for $z=2$ was first proposed by Lifshitz [1] to describe the tricritical point of liquid helium [2]. The dispersion relation (I.2) also appears in certain modified gravity scenarios involving ghost condensate [3]. In condensed matter physics, a well known example is the Rokhsar-Kivelson (RK) Quantum dimer model [4, 5] where a continuum $z=2$ Lifshitz scalar field theory emerges at the critical point.
Explicit field theoretic realization of the Lifshitz symmetry (I.1) for the special integer value of $z=2$ has been known as the quantum Lifshitz model (QLM) [6] . Various bipartite entanglement measures were subsequently analyzed within the QLM [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 16, 18, 19]. It is known that Lifshitz invariant scalar field theory admits ground state called the Rokhsar-Kivelson (RK) vacuum [4, 5], with the properties that it is local and is given by a superposition of quantum states determined by some classical probability distribution.
The point $z=2$ is however somewhat special since the Lifshitz scaling transformation (I.1) combines with the Galilean boost to form the Schrodinger group only for this value of $z$. The study of field theory beyond $z=2$ was pioneered in [20] where QLM was generalized to $(D+1)$-dimensions with $z=D$. It was also realized very interestingly that the connection to conformal field theory (CFT) found in QLM is retained for general $z=D>2$. This was further generalized in [21] where field theory for general non-integer $z$ was constructed with the help of fractional derivatives. It was found that the ground state of the Lifshitz invariant scalar field continue to take the form of RK for general $z$. Furthermore, it was found that even though the usual CFT twisted operator is not available in the Lifshitz invariant field theory, due to the special property of the RK vacuum, it is possible to employ the replica trick in a path integral formalism that they developed to study the entropic properties of these theories [21]. It was found that the results are connected to CFT for more general non-integer value of $z$. Further results in holography were also obtained in [21], and a holographic proposal for boundary Lifshitz field theory in [22].
Clearly, the physics of anisotropy is a subject of its own interest that goes beyond quantum field theory. With general applicability in mind, we explore in this paper the physical effect of anisotropy in the more basic quantum mechanics. In particular we consider time anisotropy of the background space time specified by (I.1). With the dispersion relation fixed to be of the form (I.2), quantum mechanics of free particle in the anisotropic spacetime reads
@@
where $H_{0}$ is the free Hamiltonian, $m$ is the mass of the particle and $M$ is a mass scale associated with the anisotropy. Here, defined by its action
$\partial_{j}^{z}e^{i{\bf k}\cdot{\bf x}}:=(ik_{j})^{z}e^{i{\bf k}\cdot{\bf x}}$ (I.4)
on the basis of plane wave functions, $\partial_{j}^{z}$ is an fractional derivative of order $z$, and $\Delta_{z}:=\partial_{j}^{z}\partial_{j}^{z}$ is the fractional Laplacian
$\triangle_{z}\,e^{i{\bf k}\cdot{\bf x}}=(-{\bf k}^{2})^{z/2}e^{i{\bf k}\cdot{\bf x}}.$ (I.5)
The free Hamiltonian (I.3) is Lifshitz scale invariant. One may also add a potential and consider the Lifshitz quantum mechanics defined by
@@
Note that other definitions of fractional derivatives are possible. For example, the Riesz derivative (see e.g. [24]) is defined also in terms of the Fourier transformation. However, non-analytic function $|k_{j}|$ is involved and it does not allow one to use complex contour integral theorems. For the analysis of [21], it was crucial to adopt the definition (I.4) since results such as the solution $\phi=x^{z-n}$ to the equation of motion $\partial_{x}^{z}\phi=0$ was obtained only with the help of contour deformation techniques. In this paper, we are interested in quantum mechanics that can be obtained as a low energy single particle limit of the Lifshitz QFT, therefore we consider the form (I.6) of Lifshitz quantum mechanics instead of fractional quantum mechanics defined by other types of fractional derivatives (see e.g. [24]). Note also that in general the scaling symmetry is broken except for some special form of the potential, e.g. $V\sim 1/|x|^{z}$, in which case we obtain the Lifshitz invariant quantum mechanics. In analogous to conformal quantum mechanics [23], it may be interesting to study further this kind of Lifshitz invariant quantum mechanics.
Here is a summary of the paper. In section 2, we consider the Lifshitz quantum mechanics in general. There is a belief in the literature that local conservation of probability is violated by a source term in fractional quantum mechanics. We show that this is not true for the Lifshitz quantum mechanics. In fact here the probability current is modified by anisotropy, and once the correct probability current is employed, the local conservation of probability is respected. We argue that similarly a careful identification of the probability current is needed for the other types of fractional quantum mechanics and probability is conserved in fractional quantum mechanics in general. In section 3, we consider an anisotropic Josephson Junction that is made up of an anisotropic insulating layer sandwiched between two normal superconducting materials. We show that the critical current can be significantly enhanced by anisotropy. We argue that anisotropic Josephson Junction has great potential in quantum computing.
@@ II Lifshitz Quantum Mechanics
Consider Lifshitz quantum mechanics (I.6). This kind of quantum mechanics is also refereed to as fractional quantum mechanics, see for example [24] for some discussion. It is straightforward to show that the Lifshitz quantum mechanics (I.6) is unitary and admits the local conservation law
@@
where the probability density takes the usual form $\rho=\psi^{*}\psi$, while the probability current density is modified by anisotropy. Since the fractional derivative is defined using the basis of exponential functions, $\pmb{J}$ can be most easily obtained in the momentum space and we have
@@
This result can be obtained from the Noether theorem associated with the phase symmetry $\psi\to\psi e^{i\theta}$ of quantum mechanics or by the direct manipulation of the wave equation (I.6). In fact, it is
$\psi^{*}\Delta_{z}\psi-\psi\Delta_{z}\psi^{*}=\int_{p}\int_{q}e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}),$ (II.4)
where we have denoted $\int_{p}:=\int\frac{d^{D}p}{(2\pi\hbar)^{D/2}}$. This can be identify this with a divergence term $\nabla\cdot{\bf K}$ of the form
${\bf K}=-i\int_{p}\int_{q}{\bf b}({\bf p},{\bf q})e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2})$ (II.5)
if $\bf b$} satisfies
${\bf b}\cdot({\bf q}-{\bf p})=1.$ (II.6)
Here $\bf K$} is the current density up to a normalization factor. The solution to (II.6) is not unique in 3 spatial dimensions. For example, (II.6) can be solved by (with a suitable regulator implemented, e.g. $i\varepsilon$ in the denominator),
${\bf b}=\frac{{\bf q}+{\bf p}}{{\bf q}^{2}-{\bf p}^{2}},\quad\mbox{or}\quad{\bf b}=\frac{{\bf q}-{\bf p}}{|{\bf q}-{\bf p}|^{2}},$ (II.7)
It is easy to understand the physical meaning of the non-uniqueness. Consider any two solutions ${\bf b_{1}},{\bf b_{2}}$ of (II.6), then by definition, ${\bf b_{2}}-{\bf b_{1}}$ is perpendicular to ${\bf q}-{\bf p}$. This means
${\bf b_{2}}-{\bf b_{1}}=({\bf q}-{\bf p})\times{\bf s}$ (II.8)
for some vector $\bf s({\bf p},{\bf q})$}. This then implies that the corresponding difference ${\bf K_{2}}-{\bf K_{1}}$ is given by a curl:
${\bf K_{2}}-{\bf K_{1}}=-\nabla\times\int_{p}\int_{q}{\bf s}\;e^{i({\bf q}-{\bf p})\cdot{\bf x}}\varphi^{*}({\bf p})\varphi({\bf q})((-{\bf q}^{2})^{z/2}-(-{\bf p}^{2})^{z/2}).$ (II.9)
Therefore the different solutions of (II.6) simply give rises to current densities that are different from each other by a curl term. This does not affect the expression of current conservation. For example, for the first solution in (II.7), we obtain in coordinate space,
@@
$\mathbf{J}=\lambda_{M}^{z-2}\frac{\hbar}{2mi}\frac{\nabla_{x}-\nabla_{x^{\prime}}}{\nabla_{x}^{2}-\nabla_{x^{\prime}}^{2}}\left[(\nabla_{x}^{2})^{z/2}-(\nabla_{x^{\prime}}^{2})^{z/2}\right]\psi^{*}(\mathbf{x}^{\prime},t)\psi(\mathbf{x},t)\Big|_{\mathbf{x}^{\prime}=\mathbf{x}},$ (II.10)
@@
where $\lambda_{M}:=\hbar/M$ is a length scale associated with the anisotropy. We remark that this kind of ambiguity is there even for the standard quantum mechanics although it is not usually mentioned. Like the Poynting vector, additional input from the physical setting such as symmetry, boundary condition or a good understanding of the underlying dynamics, is usually needed to fix the choice of $\pmb{J}$ properly. We remark also that the local conservation of probability in fractional quantum mechanics has been discussed in [25, 26, 27] where a source term has been found, and which has led to the conclusion of a non-conservation of probability. If this is indeed the case, it posts serious doubt in the physical consistency of the fractional quantum mechanics. However as we showed above, the probability current is indeed modified in the Lifshitz quantum mechanics and local conservation of probability is observed. We believe that the probability current considered in the fractional quantum mechanics [25, 26, 27] can also be improved similarly so as to maintain probability conservation in quantum mechanics. This point is important in order to have a consistent physical interpretation of the mathematical results computed from the fractional differential equations. In the next section, we consider Lifshitz quantum mechanics for a Josephson Junction system and predict an observable effect on the macroscopic supercurrent from the conserved probability current in the insulating layer of the Josephson Junction.
@@ III Anisotropic Josephson Junction
Suppose the Cooper pairs have typical energy $E$ in the regions 1, 2. For a junction with planar symmetry and that the insulator may be approximated by a potential barrier $V_{0}>E$ (see figure 2), the wave equation inside the insulator reads
$\lambda_{M}^{z-2}\frac{\hbar^{2}(\partial_{x}^{z})^{2}\psi}{2m}=(V_{0}-E)\psi,\quad-a<x<a,$ (III.2)
where $x$ is the direction perpendicular to the plane of the layers. Using the definition (I.4) of the fractional derivative, we obtain the general solution
@@
Here $\zeta$ is the tunneling penetration length and $\zeta_{0}$ is the tunneling penetration length in the standard Schrödinger $(z=2)$ case. Note that the penetration length is modified by anisotropy through the presence of the anisotropy length scale $\lambda_{M}$ and anisotropy index $z$. Next let us determine the tunneling current.
@@
Note that since we have assumed the Junction to be uniform in the transverse directions, the system is effectively one dimension and so the current is uniquely defined. From (II.10) we obtain
@@
The coefficients $C_{1},C_{2}$ can be determined by imposing the boundary conditions that the wavefunction matches up with the Landau-Ginzburg order parameters (III.1) of the superconductors at $x=\pm a$:
@@
Taking into account the Cooper pair has charge $q=2e$, we obtain the first Josephson equation for the tunneling current ${\mathrm{J}}:=qJ$:
${\mathrm{J}}={\mathrm{J}}_{c}\sin\varphi,\quad\varphi=\theta_{2}-\theta_{1}.$ (III.7)
Here the critical current ${\mathrm{J}}_{c}$ is given by
${\mathrm{J}}_{c}=\gamma{\mathrm{J}}_{c0}$ (III.8)
@@
${\mathrm{J}}_{c0}=\frac{q\hbar\sqrt{n_{1}n_{2}}}{m\zeta_{0}}\frac{1}{\sinh 2a/\zeta_{0}},\quad\gamma=\frac{z\zeta}{2\zeta_{0}}\frac{\sinh 2a/\zeta_{0}}{\sinh 2a/\zeta},$ (III.9)
where ${\mathrm{J}}_{c0}$ is the critical current in the isotropic case and $\gamma$ is an modification factor due to anisotropy. If we consider the application of a voltage $V$ across the Josephson Junction, then the phase becomes time dependent and we have the second Josephson equation,
@@
$\theta_{2}-\theta_{1}=(\theta_{2}-\theta_{1})_{0}+\frac{q}{\hbar c}\int_{1}^{2}{\bf A}\cdot d{\bf x},$ (III.12)
@@
Note that in the usual isotropic case, the critical current ${\mathrm{J}}_{c0}$ is exponentially suppressed by a factor of $e^{-2a/\zeta}$ since the Cooper pairs cannot tunnel deep enough into the insulating region. However this may change in the anisotropic situation. In fact, the Josephson current can be enhanced ($\gamma>1$) if the penetration length is made bigger ($\zeta>\zeta_{0}$). This happens if $z<2$ and $\lambda_{M}<\zeta_{0}$ or $z>2$ and $\lambda_{M}>\zeta_{0}$. Thus an anisotropic material has the potential of enhancing the working of Josephson Junction. In particular, if the anisotropic insulating material can be engineered such that the penetration length $\zeta$ get to become of the same order as $a$, then tunneling can become much more effective and ${\mathrm{J}}_{c}$ may be significantly enhanced. As illustration, let us consider an isotropic sample with $\zeta_{0}/2a=1/10$, in this case the critical current is suppressed with a factor of $\sim e^{-10}\sim 10^{-5}$. Introducing anisotropy, for example with $z=0.8$ and $\zeta_{0}/\lambda_{M}=10$, we can have a huge amplification factor $\gamma\sim 10^{5}$; or with $z=5$ and $\zeta_{0}/\lambda_{M}=1/10$, we have $\gamma\sim 10^{4}$. We remark that the dispersion relation (I.2) is acausal for $z<1$ due to the existence of superluminal modes [28] and the violation of the null energy conditions of the holographic dual [29]. Therefore a junction with $z\geq 1$ may be more physical. It is interesting to verify this prediction of enhanced tunneling experimentally.